I need to add thhis problem. 7 + (-9) + (-5) +6. My answer was -1. Is this correct? Please advise. Thank you.
Answer by jim_thompson5910(35256)

You are correct. 7 + (-9) + (-5) +6 = -2 + 1 = -1